### v1

# Towards Accurate and Interpretable Neuroblastoma Diagnosis via Contrastive Multi-scale Pathological Image Analysis

Abstract
Neuroblastoma, adrenal-derived, is among the most common pediatric solid malignancies, characterized by significant clinical heterogeneity. Timely and accurate pathological diagnosis from hematoxylin and eosin-stained whole slide images is critical for patient prognosis. However, current diagnostic practices primarily rely on subjective manual examination by pathologists, leading to inconsistent accuracy. Existing automated whole slide image classification methods encounter challenges such as poor interpretability, limited feature extraction capabilities, and high computational costs, restricting their practical clinical deployment. To overcome these limitations, we propose CMSwinKAN, a contrastive-learning-based multi-scale feature fusion model tailored for pathological image classification, which enhances the Swin Transformer architecture by integrating a Kernel Activation Network within its multilayer perceptron and classification head modules, significantly improving both interpretability and accuracy. By fusing multi-scale features and leveraging contrastive learning strategies, CMSwinKAN mimics clinicians’ comprehensive approach, effectively capturing global and local tissue characteristics. Additionally, we introduce a heuristic soft voting mechanism guided by clinical insights to seamlessly bridge patch-level predictions to whole slide image-level classifications. We validate CMSwinKAN on the PpNTs dataset, which was collaboratively established with our partner hospital and the publicly accessible BreakHis dataset. Results demonstrate that CMSwinKAN performs better than existing state-of-the-art pathology-specific models pre-trained on large datasets. Our source code is available at https://github.com/JSLiam94/CMSwinKAN.
{IEEEkeywords}
Neuroblastoma, Pathological Image, Classification, Kernel Activation Network, Multi-Scale Fusion
1 Introduction
[…]
Peripheral neuroblastic tumors (pNTs) are malignant tumors originating in the developing sympathetic nervous system and represent the most common solid extracranial tumors in children [1], accounting for 15% of all childhood cancer-related deaths [2, 3]. Due to the multiple subtypes of pNTs, significant differences exist in each subtype’s pathological characteristics and corresponding treatment strategies. Thus, precise subtype classification is critical for risk stratification and the development of personalized treatment plans. At present, pathological diagnosis is the gold standard for diagnosing pNTs. However, due to the subtle pathological differences between subtypes, diagnosis in clinical practice primarily relies on manual examination by pathologists, which is characterized by significant subjectivity and limitations [4, 5]. It is challenging to perform a holistic evaluation of whole slide images (WSIs). Therefore, developing automated methods to precisely identify cytopathological morphologies across entire WSIs and achieve accurate subtype diagnosis has become an urgent need.
Previous studies have developed automated WSI diagnostic methods [6, 7], yet critical challenges persist. First, the feature representation capabilities of current methods are still insufficient. Conventional machine learning approaches remain limited by handcrafted features that cannot fully characterize pathological details and tissue variations. This fundamentally constrains model performance. Second, while deep learning methods have partially overcome the limitations of handcrafted feature extraction [8], their nonlinear modeling capabilities remain constrained. Specifically, when dealing with the highly heterogeneous and complex image data of histopathology, conventional convolutional neural network (CNN) architectures struggle to efficiently capture long-range spatial dependencies and multi-scale features in the images [9]. This may further hinder the robustness of these models in practical clinical applications [10]. Moreover, the lack of model interpretability is another critical factor impeding the clinical deployment of current models [11]. Deep learning models are often considered “black boxes” and lack transparent decision-making processes. This opacity increases clinical risks and erodes pathologists’ confidence in model outputs. Such interpretability limitations significantly hinder the adoption of automated diagnostic tools in clinical practice [12].
Additionally, the size of WSI images is immense, with a single image often reaching giga pixels, far exceeding the memory capacity of a graphics processing unit (GPU) [13, 14]. As a result, the current primary solution is to divide an entire WSI image into multiple independent patches for classification. However, this patch-level processing approach poses severe challenges related to information fragmentation and false associations [15, 16]. On the one hand, the lack of practical spatial constraints between patches may result in classification outcomes that overlook the overall spatial structure of the image. On the other hand, certain critical lesion regions may be divided into separate patches and processed in isolation, impairing the comprehension of the overall pathology and leading to incomplete or inaccurate classification. These factors significantly reduce the overall diagnostic accuracy and generalization ability of the models. Therefore, there is an urgent need for a WSI classification model that can better integrate local details and global contextual relationships, possess stronger feature representation and nonlinear modeling capabilities, and offer high interpretability to facilitate the practical clinical application of automated pathology diagnostic technologies.
To address these challenges, this study proposes a novel contrastive-learning-driven multi-scale feature fusion pathological image classification model, CMSwinKAN, for classifying hematoxylin and eosin-stained(H&E) neuroblastoma (NB) histopathology patches. Additionally, a soft voting mechanism based on prior knowledge of tissue composition analysis is introduced to achieve accurate WSI classification.
Our main contributions are as follows:
  1. 1.
    Multi-scale interaction and contrastive learning: We propose a contrast-driven multi-scale feature fusion framework (CMSA) that incorporates contrastive learning and fusion across multi-scale features. This framework adaptively adjusts the contributions of features at different scales, capturing both long-range dependencies and local details in histopathological images.It resolves the insufficient interaction between multi-scale features by mitigating the negative impact of low information density in feature representation.
  2. 2.
    Lightweight and interpretable deep architecture: We introduce a multi-head self-attention module with a combined configuration of regular and shifted windows (Swin KANsformer Block) and integrate a Kernel Activation Network (KAN) as the classification head. This significantly enhances the model’s feature representation and nonlinear modeling performance. Additionally, we optimize model parameters based on feature analysis, selecting appropriate embedding dimensions and the number of neurons in KAN’s hidden layers. This achieves a lightweight design while maintaining fine-grained classification accuracy.
  3. 3.
    Soft voting mechanism for WSI-level classification: We propose a WSI classification method based on clinical prior knowledge and a soft voting mechanism. By training a support vector machine (SVM) model, we achieve fine-grained recognition of tissue components at the patch level. Combined with a dynamic weight allocation strategy, this method effectively integrates local patch features with global features, improving classification accuracy and robustness.
2 Related works
2.1 Traditional Computer Vision Models in Pathology
Traditional approaches for pathological image classification depend on handcrafted feature engineering, struggling to extract discriminative patterns from high-resolution images. For example, Gheisari et al. [17]proposed a method combining Scale-Invariant Feature Transform (SIFT) with feature encoding strategies, using an SVM classifier to classify NB pathological images into five subtypes. However, its performance was limited by the expressiveness of handcrafted features, achieving an accuracy of only 68.2%. A subsequent improvement introduced a Convolutional Deep Belief Network (CDBN) for NB pathological image classification. The CDBN model extracted high-level patterns via hierarchical feature encoding, achieving an F1 score of 86.01% under the Shimada classification system [18]. However, due to the extremely high resolution of WSIs, this approach faced challenges such as low computational efficiency, high resource requirements for whole-slide processing, and the inability to effectively model long-range spatial dependencies.
To address the high-resolution nature of WSIs, patch-based CNN methods have become mainstream. Hou et al. [19] proposed a decision fusion model that improved recognition performance by integrating local patch classification results. However, its ”independent processing-simple fusion” approach led to fragmented spatial information. Recently, Li et al. [20] proposed a weakly supervised method based on the Variational Information Bottleneck, further refining the optimization of feature representation. While this method performed well on multiple datasets, its patch-level processing still fell short in capturing global structural information effectively.
2.2 Development of Pathology Large Models
Recent advances in self-supervised learning (SSL) have propelled breakthroughs in large pathology models. For instance, Wang et al. proposed the CTransPath model, which combines CNNs with a multi-scale Swin Transformer architecture, achieving state-of-the-art (SOTA) performance across multiple pathology tasks [21]. Additionally, Chen et al. introduced the universal pathology pretraining model UNI, which was pre-trained on over 1 million pathological images. UNI surpassed existing methods across 34 pathology tasks, achieving comprehensive breakthroughs from tissue classification to cancer subtype prediction [22].
The introduction of vision-language models has further enhanced the generalization capabilities of large pathology models. Lu et al. proposed the CONCH model, which was pre-trained on 1.17 million image-text pairs in a task-agnostic manner, achieving leading performance in tasks such as pathological image classification, segmentation, and retrieval [23]. Furthermore, Azizi et al. developed the REMEDIS framework, which improved the robustness of medical image models through SSL, enabling high performance even in out-of-distribution scenarios [24].
Figure 1: Overall architecture of the CMSwinKAN model. The model consists mainly of the Swin KANsformer block and the contrast-driven Multi-Scale Feature Fusion Module. Within the Swin KANsformer block, W-MSA and SW-MSA are multi-head self-attention modules with regular and shifted window configurations, respectively.
2.3 Advantages of Lightweight Small Models
Despite strong performance, pathology large models face four key limitations: high computational resource consumption (e.g., months of GPU pretraining), limited deployability on medical edge devices, lack of interpretability (due to complex model structures that are difficult to map to clinically interpretable pathological features), and poor adaptability to small-sample datasets, such as pediatric tumors.
In contrast, lightweight, small models offer unique advantages in terms of computational efficiency and adaptability. For example, Zormpas-Petridis et al. proposed the SuperHistopath framework, which uses the Simple Linear Iterative Clustering superpixel algorithm to segment low-resolution WSIs and classify superpixels based on CNNs, enabling precise mapping of tumor heterogeneity [25]. Ding et al. introduced the Fuzzy-Guided Multi-Granularity Deep Neural Network (FMDNN), which simulates the multi-level diagnostic process of pathologists and outperformed existing classification methods on several public datasets [26].
In the field of efficient Transformer design, Stegmüller et al. proposed ScoreNet, which uses a differentiable recommendation mechanism to focus on the most discriminative pathological regions and incorporates the ScoreMix data augmentation strategy to enhance generalization [27]. Additionally, Liu et al. developed the DetexNet network, which combines Texture Energy Measure to improve the representation ability of low-level pathological features, enabling high accuracy on small-sample pathology datasets [28]. Although these methods are significantly more efficient than large models, they still lack sufficient capacity to model local features such as nuclear details and tissue structure patterns while effectively integrating multi-scale information.
2.4 Multiscale Methods in Medical Image Processing
Multiscale feature extraction and fusion have long been critical for improving the performance of image processing models. In the field of natural images, Chen et al. proposed the CrossViT model, which uses a dual-branch Transformer to process image patches of different sizes and fuses them via a cross-attention mechanism, effectively improving image classification [29]. Li et al. introduced MViTv2, which improved multiscale visual Transformers by decomposing relative position encoding and incorporating residual pooling connections, achieving efficient performance in classification and detection tasks [30]. Ren et al. proposed Shunted Self-Attention, which includes heterogeneous receptive fields at each layer to aggregate multiscale features of tokens, effectively addressing the limitations of traditional self-attention in handling multiscale objects [31]. Additionally, Jiao et al. introduced DilateFormer, which uses multiscale dilated attention to expand the receptive field while maintaining computational efficiency, achieving SOTA performance across multiple vision tasks [32].
In the medical image domain, Huo et al. proposed the Hierarchical Multiscale Feature Fusion Network (HiFuse), which integrates global and local feature blocks and uses adaptive feature fusion modules to combine information from different scales efficiently, significantly enhancing model robustness in complex medical image classification tasks [33]. Chen et al. proposed an accurate leukocyte detection model based on deformable-DETR integrated with multi-level feature fusion, effectively capturing detailed cellular characteristics through deformable self-attention and hierarchical fusion mechanisms, thus significantly enhancing leukocyte detection accuracy [34]. Lei et al. introduced the ConDSeg framework, which uses a contrast-driven feature enhancement strategy to improve boundary recognition in image segmentation tasks, achieving SOTA performance on multiple pathology datasets [35]. Given its outstanding performance in medical image segmentation, we incorporate similar strategies into pediatric pathology image classification to address challenges such as multiscale information fusion and boundary ambiguity, thereby improving classification accuracy and model generalization capabilities.
3 Method
3.1 Overall Architecture
Fig. 1 illustrates the overall architecture of the CMSwinKAN model (micro version), which is designed to address the challenges of multi-scale feature representation and spatial dependency modeling in pathological image classification. This model is specifically tailored for the patch-level classification of pathological images from H&E-stained slides by combining a lightweight architecture with clinical interpretability. The model consists of three core modules:
1). Swin KANsformer Block: This is an improved computational unit based on the Swin Transformer, where the traditional multi-layer perceptron (MLP) layers are replaced with KAN layers. This substitution effectively boosts both local feature representation and interpretability.
2). CMSA Module: By leveraging contrastive learning and an adaptive feature weighting mechanism, this module integrates multi-scale pathological patterns (e.g., nuclear details and tissue structures), thereby simulating the comprehensive diagnostic strategies of clinical pathologists.
3). KAN Classification Head: The KAN layers replace the conventional fully connected layers in the classification head, improving the model’s non-linear representation capacity while reducing computational complexity.
With the introduction of KAN layers, the model significantly reduces parameter count by utilizing smaller embedding dimensions (e.g., 24 for the micro version and 56 for the tiny version, compared to 128 for Swin-base). This reduction comes with minimal performance degradation; in fact, the feature representation and generalization capabilities are enhanced. Furthermore, the Swin KANsformer Block alternates between Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA) layers, replacing all MLP layers with KAN layers, which further improves the model’s efficiency and interpretability.
The core structure of Contrastive-Driven Feature Augmentation (CDFA) in the CMSA module enhances multi-scale feature fusion by comparing attention across different scales, resulting in richer and more diagnostically meaningful insights. The three modules work synergistically, enabling the CMSwinKAN model to efficiently and accurately classify H&E-stained pathological image patches and laying the foundation for whole-slide image classification tasks.
3.2 Swin KANsformer Block
The Swin Transformer Block has demonstrated numerous advantages in image processing and medical image analysis compared to traditional Vision Transformers. Introducing hierarchical feature maps and a local window-based self-attention mechanism effectively reduces computational complexity while enabling the extraction of multi-scale features, making it particularly well-suited for handling high-resolution images. Additionally, its shifted window mechanism enhances information exchange between windows, improving the model’s representation and feature extraction capabilities. However, the Swin Transformer Block has limitations, such as constrained global receptive fields, high computational complexity, and reliance on large-scale pretraining datasets. To address these shortcomings, we present the Swin KANsformer Block, an optimized enhancement of the original Swin Transformer Block.
The Swin KANsformer Block replaces the traditional MLP layers in the Swin Transformer Block with KAN layers, addressing the deficiencies of the Swin Transformer when processing complex H&E-stained pathological image data. Inheriting the structure of the Swin Transformer Block, its multi-head self-attention layers alternate between two types in consecutive Swin KANsformer Blocks: W-MSA layers and SW-MSA layers. Within the model, the input RGB H&E-stained pathological images are first divided into non-overlapping 4×4 image patches via the Patch Partition module, with each patch having a feature dimension of 4×4×3 = 48. A linear embedding layer then projects these patches into an arbitrary feature dimension, denoted as C.
To facilitate the preservation of multi-scale features, we redefine the model’s stages. Aside from the final stage, which only contains $d_{3}$ Swin KANsformer Blocks, each stage includes $d_{i}$ Swin KANsformer Blocks and a patch merging layer. In our base model, the number of Swin KANsformer Blocks across different stages is defined as $\left[d_{0},d_{1},d_{2},d_{3}\right]=[2,2,8,2]$. Due to the Patch Merging layers, the spatial dimensions of image features are progressively reduced after the first three stages, with sizes $\left[\left(\frac{H}{8},\frac{W}{8}\right),\left(\frac{H}{16},\frac{W}{16}\right),\left(\frac{H}{32},\frac{W}{32}\right)\right]$.
3.3 Contrastive-Driven Multi-Scale Aggregation Module
Analyzing images at multiple scales leverages scale-aware knowledge to enhance model performance and uses the learned cross-scale relationships to maintain spatial consistency. We employ the CMSA module to utilize multi-scale feature information effectively. This module introduces the CDFA module, which addresses two significant challenges in medical image analysis: “soft boundaries” and co-occurrence phenomena. In our model, the CDFA module serves as the fusion mechanism for multi-scale features extracted from different layers of the Swin KANsformer.
Features from the first three stages of the Swin KANsformer are stored in a feature matrix and multiplied by a set of learnable weights $\left[i,j,k\right]$ to dynamically adjust the contribution of each stage. The features from the 2nd and 3rd stages, with sizes $[\left(\frac{H}{16},\frac{W}{16}\right),\left(\frac{H}{32},\frac{W}{32}\right)]$, are upsampled via bilinear interpolation to the size of the first stage $\left(\frac{H}{8},\frac{W}{8}\right)$. The feature channels $[2C,4C,8C]$ are unified to $2^{s}C$ using 1×1 convolution blocks (in the base model, s is set to 1, so the features from the 1st stage do not require convolutional processing). After these operations, the features are fed into the CDFA module.
In the CDFA module, at each spatial location $(i,j)$, attention weights within a $K\times K$ window centered at $(i,j)$ are computed by combining details from shallow and deep features. Specifically, the processed shallow features (features retained from the 1st and 2nd stages) $F_{1}$ and $F_{2}$ are passed through two different linear layers to compute attention weights $A_{1}$ and $A_{2}$, respectively, where $A_{1},A_{2}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times K^{2}}$. The attention weight computation can be expressed as:
$A_{1}=W_{1}\cdot F_{1},A_{2}=W_{2}\cdot F_{2},$ (1)
where $W_{1},\ W_{2}\in\mathbb{R}^{2C\times K^{4}}$ are the linear transformation weight layers for the multi-scale feature maps from the 1st and 2nd stages, respectively. Each spatial location $(i,j)$ in the two attention maps is then reshaped into ${\hat{A}}_{1_{i,j}}\in\mathbb{R}^{K\times K\times K^{2}}$ and ${\hat{A}}_{2_{i,j}}\in\mathbb{R}^{K\times K\times K^{2}}$, both activated by softmax.
The processed deep feature (retained from the 3rd stage) $F_{3}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 2C}$ is initially fused through two CBR blocks. A CBR block combines a convolution layer, batch normalization, and a ReLU activation function, simplifying the network construction process and improving training efficiency and performance. The fused 2C-dimensional feature vector is then projected through a linear layer to a value vector $\bm{V}_{3}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 2C}$. This value vector $\bm{V}_{3}$ is then unfolded over local windows, aggregating neighborhood information for each position. Let $\bm{V}_{\Delta_{i,j}}\in{R^{2C\times{K^{2}}}}$ represent all values within a local window centered at $\left(i,j\right)$, defined as:
$\bm{V}_{\Delta_{i,j}}=\{\bm{V}_{i+p-\left\lfloor\frac{K}{2}\right\rfloor,j+q-\left\lfloor\frac{K}{2}\right\rfloor}\},\qquad 0\leq p,q<K.$ (2)
The aggregation is performed in two steps:
$\hat{\bm{V}}_{\Delta_{i,j}}=Softmax(\hat{A}_{1_{i,j}})\otimes(Softmax(\hat{A}_{2_{i,j}})\otimes(\bm{V}_{\Delta_{i,j}}),$ (3)
where $\otimes$ denotes matrix multiplication. Finally, the weighted feature values are densely aggregated to produce the final output feature map. Specifically, the aggregated feature at position $(i,j)$ is:
${\widetilde{F}}_{i,j}=\sum_{0\leq=m,n<K}{\widetilde{\bm{V}}}_{\triangle_{i+m-\left\lfloor\frac{K}{2}\right\rfloor,j+n-\left\lfloor\frac{K}{2}\right\rfloor}}^{i,j}.$ (4)
3.4 KAN Classification Head
The multi-scale fused features obtained from the CMSA module are concatenated with the output residuals from the 4th stage of the Swin KANsformer. This step appropriately preserves the detail and semantic information from the original image, enabling the multi-scale fused features to guide image classification effectively. We employ the KAN as the classification head to endow the model with strong non-linear fitting capabilities and superior classification performance. The KAN processes input features and outputs classification probabilities. Each activation function in KAN is a learnable spline function, which can be expressed as:
$\phi(x)=w_{b}b(x)+w_{s}spline(x),$ (5)
where $b(x)$ is a fixed basis function (e.g., SiLU), and $spline(x)$ is a learnable spline function.
KAN uses B-spline basis functions to construct piecewise polynomials with local support (i.e., each basis function is non-zero only within a finite interval). This property allows the model to handle high-dimensional data efficiently.
The B-spline basis function serves as the core component of KAN for constructing piecewise polynomials. For input data $x$ and a given set of grid points $t=\{t_{0},t_{1},\ldots,t_{n}\}$, the B-spline basis function $N_{i,k}\left(x\right)$ is computed as follows:
$N_{i,k}\left(x\right)=w_{i,k}\left(x\right)\cdot N_{i,k-1}\left(x\right)+(1-w_{i+1,k}(x))\cdot N_{i+1,k-1}(x),$ (6)
where
$w_{i,k}\left(x\right)=\frac{x-t_{i}}{t_{i+k}-t_{i}}.$ (7)
The recursive formulation $N_{i,k}\left(x\right)$ ensures efficient computation of B-spline basis functions, while their local support lets the KAN head model complex feature spaces efficiently.
3.5 WSI Voting Classification Mechanism
Figure 2: The architecture of the proposed WSI voting mechanism. The original pathological image is standardized, cut, and stacked to generate a sequence of slices. After feature classification of the slices using an SVM classifier, weights are dynamically assigned based on classification confidence, ultimately implementing a weighted soft voting strategy for decision.
To accurately classify the entire WSI and assist in pathological diagnosis, we design a heuristic soft voting classification process based on clinical knowledge rules, as illustrated in Fig. 2. This process is grounded in the classification results of individual patches and incorporates global feature information to ultimately determine the overall classification label of the WSI. Unlike traditional hard voting methods, the soft voting mechanism we employ flexibly assigns weights based on the varying importance of slices, yielding more robust classification. Furthermore, to more finely discern specific tissue structures within slices, such as neuropil or Schwannian stroma, we train an SVM model to identify tissue components accurately. Below, we will elaborate on the voting mechanism’s design and the SVM model’s training and application.
3.5.1 Voting Mechanism Design
The implementation of WSI classification through a hard voting mechanism is straightforward; it treats the classification results of all patches as equally weighted votes, determining the final classification based on the majority rule. However, the precision of the hard voting method is relatively low. To achieve a more precise classification, we incorporated prior knowledge and designed a soft voting approach for WSI classification. According to principles of pathological analysis, if a patch contains regions with neuropil, it suggests the presence of cell nests, indicating that this patch is likely to belong to a malignant category, specifically ganglioneuroblastoma (GNB) or NB. Conversely, if a patch contains regions of Schwannian stroma, it suggests the absence of cell nests, leaning more towards ganglioneuroma (GN). Furthermore, if Schwannian stroma constitutes 50% or more of the patch, the WSI tends towards GNB or GN; otherwise, it leans towards NB.
Based on this knowledge, we can derive the voting weights for all patches within the WSI, treating each patch’s weight as a displacement vector $\vec{w}$ in the classification space, where the direction corresponds to the category set and the length represents the weight size. By integrating the voting weights of all patches with their classification results, we arrive at the final WSI classification ${\mathrm{Label}}_{\mathrm{WSI}}$:
$P\in RN\times C,\qquad P\in R^{N\times C},$
$W\in RN\times C,\qquad W\in R^{N\times C},$
$\mathbf{s}=(W\circ P)^{\top}\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{N}w_{i,1}p_{i,1}\\ \sum\limits_{i=1}^{N}w_{i,2}p_{i,2}\\ \vdots\\ \sum\limits_{i=1}^{N}w_{i,C}p_{i,C}\end{bmatrix},$ (8)
$\alpha=\begin{bmatrix}1&1&\cdots&1\end{bmatrix}\left(W\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix}\right)=\sum\limits_{i=1}^{N}\sum\limits_{c=1}^{C}w_{i,c},$ (9)
$\hat{\mathbf{p}}=\frac{\mathbf{s}}{\alpha}=\frac{1}{\alpha}\begin{bmatrix}\sum\limits_{i=1}^{N}w_{i,1}p_{i,1}\\ \sum\limits_{i=1}^{N}w_{i,2}p_{i,2}\\ \vdots\\ \sum\limits_{i=1}^{N}w_{i,C}p_{i,C}\end{bmatrix},$ (10)
$\mathrm{Label}_{\mathrm{WSI}}={\arg\max}\ \hat{p}_{c},\qquad c\in\{1,\ldots,C\},$ (11)
where $P_{\mathrm{i},C}$ represents the classification category and probability of the C-th patch, and $w_{\mathrm{i},C}$ denotes the corresponding voting weight.
The soft voting mechanism integrates patch classification results, facilitating the transfer of classification information from local to global levels. This weighted mechanism better accommodates the high heterogeneity issues in pathological image classification. Additionally, it simplifies the final steps of WSI classification by converting complex probability distributions into clear classification labels, providing a straightforward diagnostic basis for practical applications.
3.5.2 Training and Application of the SVM Model
Leveraging the features extracted by the patch classification model CMSwinKAN, we train a multi-class SVM model to determine whether a patch contains neuropil, Schwannian stroma, or other tissue components. During the inference phase, the SVM model receives the input feature x of each patch and outputs the corresponding category probability distribution $P_{\mathrm{SVM}}$:
$P_{\text{SVM}}=\text{softmax}\left(\begin{array}[]{c}\mathbf{w}_{0}\cdot\bm{\phi}(\mathbf{x})+b_{0},\\ \mathbf{w}_{1}\cdot\bm{\phi}(\mathbf{x})+b_{1},\\ \mathbf{w}_{2}\cdot\bm{\phi}(\mathbf{x})+b_{2}\end{array}\right),$ (12)
where the “softmax” function normalizes the SVM output values into a probability distribution. $P_{\mathrm{SVM}}=\left[P_{0},P_{1},P_{2}\right]$ represents the probabilities of the patch belonging to other tissues, neuropil, and Schwannian stroma, respectively. Based on the SVM prediction results, the voting weight of the patch $w_{\mathrm{patch}}$ is determined as follows:
$w_{\mathrm{patch}}=\alpha\cdot P_{1}+\beta\cdot P_{2},$ (13)
where $P_{1}$ and $P_{2}$ are the neuropil and Schwannian stroma probabilities, respectively. $\alpha$ and $\beta$ are weighting coefficients used to control the importance of different tissue components, with $\alpha>\ \beta$ emphasizing the significance of neuropil. If the SVM classifies the patch as other tissue, then $w_{\mathrm{patch}}=\gamma$, where $\gamma$ is a lower baseline weight.
Figure 3: Medical data processing workflow based on pathological sections, which includes four main stages: (1) Data Collection, covering the collection of clinical patient tissue samples, tissue sample processing, and staining, as well as WSI scanning; (2) WSI preprocessing, involving physician evaluation and screening of WSIs and data organization; (3) Patch-level Data Construction, including determining patch size and sliding window segmentation of slices; (4) Dataset Preprocessing, covering invalid region filtering and dataset splitting, ultimately providing structured data support for model training and evaluation.
Table 1: Different histological types and gender composition of our private pNTs dataset.
Histological Type Total Cases (Percentage) WSI Male Female
Undifferentiated (UD) 20 (19.6%) 63 14 6
Poorly Differentiated (PD) 40 (39.2%) 56 24 16
Differentiated (D) 22 (21.6%) 62 8 14
Ganglioneuroblastoma (GNB) 10 (9.8%) 63 4 6
Ganglioneuroma (GN) 10 (9.8%) 60 3 7
Total 102 (100.0%) 304 53 49
4 Experiments
4.1 Data Collection and Processing
The academic community faces a scarcity of large-scale, fine-grained datasets for NB due to the complexity and resource-intensive nature of data collection and annotation. This limitation restricts the availability of high-quality resources for research and development in this domain. To address this gap, we collaborated with the Department of Pediatric Oncology at Zhejiang University Children’s Hospital to establish a private dataset, Pediatric peripheral Neuroblastic Tumors(PpNTs). As shown in Table 1, this dataset includes 102 children diagnosed with NB at Zhejiang University Children’s Hospital from January 2019 to December 2023, comprising a total of 304 WSIs. Digitization was performed using a Pannoramic SCAN II digital scanner (3DHISTECH Ltd, Budapest, Hungary) at a magnification of 20×20$\times$20 ×.
4.1.1 Sliding Window Segmentation
As shown in Fig. 3, a sliding window algorithm is employed with a non-overlapping strategy. The coordinates start from the top-left corner of the WSI (0,0), and the window slides horizontally with a step size of 512 pixels. We discard these parts for the right and bottom areas that cannot be entirely covered by a $512\times 512$ pixel window, which are primarily blank regions, to avoid introducing invalid information.
4.1.2 Generation of Image Patches
Following the slicing of WSI, numerous image patches are generated, including non-tissue regions such as blank backgrounds and areas densely populated with red blood cells. These non-informative patches may introduce noise and compromise the model’s ability to effectively extract discriminative features.
4.1.3 Data Quality Improvement
To enhance data quality, this study adopts a two-stage filtering rule: White Pixel Filtering: The proportion of white pixels (RGB value $\geq$ 245) in a patch is calculated; if it exceeds 50%, the patch is deemed an invalid area such as blank background or staining defect and is discarded. Red Blood Cell Area Filtering: If the proportion of red pixels exceeds 50%, the patch is classified as a dense red blood cell area, which may obscure key pathological features and is therefore discarded.
Figure 4: Visualization of our private PpNTs dataset, which contains five types of neuroblastic tumors: Ganglioneuroma, Ganglioneuroblastoma, Poorly Differentiated Neuroblastoma, Undifferentiated Neuroblastoma, and Differentiated Neuroblastoma.
Table 2: Number of WSIs and patches per category in our private PpNTs dataset.
GN GNB PD D UD
Class WSI Patch WSI Patch WSI Patch WSI Patch WSI Patch
Train 48 $536{,}013$ 50 $484{,}792$ 44 $432{,}627$ 49 $474{,}440$ 50 $451{,}986$
Test 12 $110{,}519$ 13 $138{,}995$ 12 $116{,}903$ 13 $151{,}007$ 13 $103{,}438$
Table 3: Settings and parameters of different versions of CMSwinKAN models.
Model embed_dim num_heads Params FLOPs
CMSwinKAN-mini 24 (2, 4, 8, 16) 1.29M 0.21G
CMSwinKAN-micro 24 (4, 8, 16, 32) 1.32M 0.22G
CMSwinKAN-tiny 56 (4, 8, 16, 32) 5.52M 1.08G
4.1.4 Dataset Construction
After rigorous screening, the training and testing sets are divided based on the pathological category labels of the WSIs in an $80:20$ ratio. This results in a training set containing $2{,}379{,}858$ patches and a testing set containing $620{,}862$ patches, ensuring consistent class distribution across both datasets to avoid model performance bias due to sampling variance. Each WSI from the same case appears in only one subset to prevent data leakage. Fig. 4 provides a visual overview of our dataset, while Table 2 presents the number of WSIs and patches for each category.
4.2 Comparison on Existing Datasets
To validate the model’s generalization ability, we used the public BreakHis dataset for further testing. The BreakHis dataset is a public dataset for classifying breast cancer histopathological images collected by the P&D laboratory in Brazil. It consists of $7{,}909$ histopathological biopsy images from $82$ patients, categorized into different magnifications (×$\times 40$, ×$\times 100$, ×$\times 200$, ×$\times 400$) across a total of $8$ categories: $4$ benign types (adenosis,fibroadenoma,phyllodes tumor,tubular adenoma) and $4$ malignant types (ductal carcinoma,lobular carcinoma,mucinous carcinoma,papillary carcinoma). To thoroughly assess the generalization capability of our model(which differs the most to our own dataset in magnification), we selected images at a magnification of ×$\times 400$ and utilized the $8$ subcategories as the experimental dataset. The dataset was split into training and testing sets at an $80:20$ ratio, resulting in a certain number of WSIs. To address the class imbalance, we performed data augmentation on the training set, including brightness, contrast, and HSV color transformations, increasing the original $1{,}427$ patches to $6{,}109$ patches, while the testing set remained at $393$ patches.
4.3 Experimental Setup
The experiments were conducted on an NVIDIA Tesla V100-SXM2 GPU (32GB graphic memory) hardware platform with a software environment based on Python 3.8.20 and PyTorch 2.4.1 with CUDA 12.1 acceleration library. The training process on our private dataset lasted for 30 epochs, with a batch size of 64, taking approximately 75 hours.
For training data, input images were resized to $224\times 224$ pixels, and data augmentation techniques, such as random horizontal flipping, were applied to enhance robustness. Images were directly resized to $224\times 224$ pixels for the testing data. The model was not pre-trained; instead, model parameters were randomly initialized using a He normal distribution[36], and the AdamW optimizer was employed in conjunction with a cosine annealing algorithm. A cosine annealing strategy was utilized during training, and gradient accumulation was employed with a step count of four, effectively expanding the global batch size to 256, enhancing computational efficiency and numerical stability through AMP-mixed precision training. The model optimized for the cross-entropy loss, with a fixed batch size of 64, and implemented 20% structured pruning on the convolutional layers to reduce computational complexity.
We set $\alpha=1$ for neuropil and $\beta=8$ for Schwannian stroma, with $\gamma=1$ assigned to non-key tissue regions. This weighted fusion approach enhances diagnostic specificity by emphasizing critical tissue features while maintaining balanced contributions from all regions.
Table 4: Comparison of CMSwinKAN with other advanced models on the private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC (%)
General methods
VGGNet [37] 81.89 81.52 74.37 81.76 97.02
ResNet-50 [38] 92.75 92.35 90.94 92.73 99.43
DenseNet [39] 92.02 91.73 89.71 91.93 99.36
Vision Transformer [40] 89.24 89.12 85.90 89.23 98.79
Swin Transformer [41] 91.81 91.68 89.64 91.81 99.30
Pathological large-scale pre-trained general models
UNI [22] 93.35 93.12 91.64 93.29 99.43
CTransPath [21] 92.22 92.11 90.22 92.24 99.36
CONCH [23] 92.17 92.01 90.32 92.15 95.66
Pathological classification model
HisImage [42] 86.29 86.26 82.78 86.21 98.00
FMDNN [26] 92.53 92.30 89.81 92.48 99.41
StoHisNet [43] 92.21 92.03 90.87 92.18 99.33
HiFuse [33] 91.49 91.40 89.31 91.49 99.29
Ours
CMSwinKAN-micro 93.58 93.42 92.12 93.57 99.53
Figure 5: Confusion matrixes of CMSwinKAN on our private PpNTs and public BreakHis tested datasets.
4.4 Evaluation Metrics
To comprehensively evaluate the performance of different models across various datasets, we employed the following standard evaluation metrics: Accuracy (ACC), Balanced Accuracy (BACC), Weighted F1 Score, Cohen’s Kappa Coefficient, and Area Under the Curve (AUC). The calculation of BACC requires first computing Recall and Precision.
$ACC=\frac{TP+TN}{TP+TN+FP+FN}\\ ,$ (14)
$BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)\\ ,$ (15)
$Precision=\frac{TP}{TP+FP}\\ ,$ (16)
$Recall=\frac{TP}{TP+FN}\\ ,$ (17)
$F1=2\times\frac{{Precision}\times{Recall}}{{Precision}+{Recall}}\\ ,$ (18)
$kappa=\frac{p_{0}-p_{e}}{1-p_{e}}\\ ,$ (19)
$AUC=\frac{\sum(pi,nj)_{pi>nj}}{P\times N}.$ (20)
Figure 6: ROC curves of CMSwinKAN on our private PpNTs and public BreakHis tested datasets.
In the above formulas, True Positive $(TP)$ and True Negative $(TN)$ represent the samples correctly predicted as positive and negative, respectively, while False Positive $(FP)$ and False Negative $(FN)$ represent the samples incorrectly predicted as positive and negative, respectively. $P$ denotes the number of positive samples, and $N$ denotes the number of negative samples. $p_{i}$ represents the predicted score for positive samples, indicating the probability of predicting a positive sample as a positive case; $n_{j}$ represents the predicted score for negative samples, indicating the probability of predicting a negative sample as a positive case.
Table 5: Comparison of CMSwinKAN with other advanced models for WSI classification on the private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%)
General methods
VGGNet [37] (hard vote) 92.06 91.79 90.07 91.88
VGGNet [37] (soft vote) 93.65 93.46 92.06 93.55
ResNet-50 [38] (hard vote) 98.41 98.33 98.01 98.39
ResNet-50 [38] (soft vote) 98.41 98.33 98.01 98.39
DenseNet [39] (hard vote) 98.41 98.33 98.01 98.39
DenseNet [39] (soft vote) 98.41 98.33 98.01 98.39
Vision Transformer [40] (hard vote) 98.41 98.33 98.01 98.39
Vision Transformer [40] (soft vote) 98.41 98.33 98.01 98.39
Swin Transformer [41] (hard vote) 98.41 98.33 98.01 98.39
Swin Transformer [41] (soft vote) 98.41 98.33 98.01 98.39
Pathological large-scale pre-trained general models
UNI [22] (hard vote) 98.41 98.33 98.01 98.39
UNI [22] (soft vote) 98.41 98.33 98.01 98.39
CTransPath [21] (hard vote) 98.41 98.33 98.01 98.39
CTransPath [21] (soft vote) 98.41 98.33 98.01 98.39
CONCH [23] (hard vote) 98.41 98.33 98.01 98.39
CONCH [23] (soft vote) 98.41 98.33 98.01 98.39
Pathological classification models
HisImage [42] (hard vote) 95.24 95.13 94.04 95.20
HisImage [42] (soft vote) 96.83 96.79 96.03 96.82
FMDNN [26] (hard vote) 98.41 98.33 98.01 98.39
FMDNN [26] (soft vote) 98.41 98.33 98.01 98.39
StoHisNet [43] (hard vote) 98.41 98.33 98.01 98.39
StoHisNet [43] (soft vote) 98.41 98.33 98.01 98.39
HiFuse [33] (hard vote) 98.41 98.33 98.01 98.39
HiFuse [33] (soft vote) 98.41 98.33 98.01 98.39
Ours
CMSwinKAN-micro (hard vote) 98.41 98.33 98.01 98.39
CMSwinKAN-micro (soft vote) 98.41 98.33 98.01 98.39
4.5 Comparative Experiments
4.5.1 Model Performance Comparison
Table 4 shows that we conducted comparative experiments on the PpNTs dataset, evaluating our model against other typical classification models, general large models in pathology, and advanced pathological image classification models. We validated the effectiveness of the models in both microscopic image classification and WSI classification. Compared to other typical models, general large models in pathology, or advanced pathological image classification models, CMSwinKAN demonstrates significant advantages in handling microscopic image classification tasks. Fig. 5 and Fig. 6 demonstrate the confusion matrices and AUC curves of CMSwinKAN tested on the PpNTs and BreakHis datasets.
Compared to the baseline models Swin Transformer and Vision Transformer, the ACC is improved by 1.77% and 4.34%, respectively. This improvement can be attributed to integrating the CMSA module and KAN in CMSwinKAN, which effectively captures fine-grained features and high-level semantic information in pathological images. This capability compensates for the limitations of the Swin Transformer in
Figure 7: Relationship between ACC, parameters, and FLOPs of CMSwinKAN and other advanced models on our private PpNTs dataset. The size of each circle represents the number of parameters for each model.
terms of field of view and lack of contextual information interaction, as well as the Vision Transformer’s shortcomings in capturing local features.
Compared with general large models in pathology, our model, which uses random initialization without pre-training, still outperformed CONCH and UNI, which were pre-trained on large-scale pathological datasets and transferred to our private dataset. While these models perform well in multimodal pathological data, their effectiveness in capturing subtle features and differentiating similar cell images in specialized fields, such as NB image classification, is limited. As illustrated in Fig. 7, CMSwinKAN demonstrates a distinct advantage in combining high accuracy with a lightweight design compared to other models.
Additionally, we compared our CMSwinKAN with advanced pathological image classification models, such as HisImage [42], FMDNN [26], and HiFuse [33], which have limitations in handling similar categorical features. Specifically, under the same training conditions, we surpassed them in all metrics; for instance, their KAPPA scores were 82.78%, 89.81%, and 89.31%, respectively, while our CMSwinKAN model achieved a KAPPA score of 92.12%. In terms of BACC, CMSwinKAN increased performance by 7.16%, 1.12%, and 2.02% over these models, respectively, significantly enhancing pathological diagnostic capabilities. Compared to the improved baseline model CTransPath based on Swin Transformer and StoHisNet using Swin Transformer Block, our method improved the ACC by 1.37%, demonstrating the architectural advantages of CMSwinKAN. In the context of WSI classification, in Table 5, the CMSwinKAN model exhibits clinical application value. After adopting a soft voting strategy, it achieved excellent performance with 98.41% ACC and 98.33% BACC, comparable to mainstream architectures like ResNet-50, DenseNet, and large pathological models like UNI. Notably, this strategy performed exceptionally well in baseline models: VGGNet’s ACC improved from 92.06% to 93.65% and BACC from 91.79% to 93.46%, confirming the beneficial impact of weighted fusion on diagnostic decisions. The model innovatively reduced the parameter count to 1.53% of that of the Swin Transformer through structural optimization, achieving an enhancement in parameter utilization efficiency while maintaining comparable WSI classification performance (98.41% ACC). This high-performance and high-efficiency characteristic allows CMSwinKAN to meet the fine-grained analysis needs of microscopic images with a KAPPA score of 92.12% and a maximum BACC improvement of 6.6% while adapting to large-scale clinical WSI processing scenarios, providing a new technical paradigm for intelligent pathological diagnosis.
Figure 8: The loss-gradient plots for different versions of the CMSwinKAN models evaluated on our private PpNTs and public BreakHis datasets.
4.5.2 Validation of Model Generalization Ability
To further validate the generalization capability of the proposed CMSwinKAN model, we compared it with other typical classification models, large models in pathology, and advanced pathological image classification models on the publicly available breast cancer dataset, BreakHis. Table 6 shows that CMSwinKAN demonstrated strong generalization ability on the imbalanced dataset, achieving excellent results across all metrics. Fig. 8 depicts the loss and gradients on the test set during training.
Table 6: Comparison of CMSwinKAN with other advanced models on the public BreakHis dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC (%)
General methods
VGGNet [37] 48.09 23.57 21.44 41.37 67.71
ResNet-50 [38] 68.45 55.61 47.65 67.40 92.37
DenseNet [39] 83.21 80.81 77.33 83.64 98.25
Vision Transformer [40] 68.45 59.84 61.27 67.58 91.03
Swin Transformer [41] 75.82 71.22 68.38 76.06 96.37
Pathological large-scale pre-trained general models
UNI [22] 78.88 69.91 72.07 71.41 97.03
CTransPath [21] 83.46 86.62 78.92 83.79 98.52
CONCH [23] 86.77 86.99 86.98 86.81 94.73
Pathological classification models
HisImage [42] 64.63 55.96 52.50 57.59 84.43
FMDNN [26] 86.32 81.68 84.71 86.14 96.57
StoHisNet [43] 74.84 68.33 67.32 74.24 94.26
HiFuse [33] 62.84 47.93 48.78 60.59 87.02
Ours
CMSwinKAN-tiny 87.79 87.56 85.69 88.04 98.80
Table 7: Comparison of CMSwinKAN with other advanced models using hard voting for WSI classification (BreakHis).
Model ACC (%) BACC (%) KAPPA (%) F1 (%)
General methods
VGGNet [37] (hard vote) 46.91 33.45 32.00 23.60
ResNet-50 [38] (hard vote) 60.49 63.36 51.98 54.98
DenseNet [39] (hard vote) 93.83 94.21 91.75 92.67
Vision Transformer [40] (hard vote) 66.67 60.28 55.50 57.88
Swin Transformer [41] (hard vote) 80.25 75.54 73.80 73.71
Pathological large-scale pre-trained general models
UNI [22] (hard vote) 88.89 78.99 84.70 82.49
CTransPath [21] (hard vote) 92.59 94.56 90.17 91.84
CONCH [23] (hard vote) 95.05 93.75 93.28 94.51
Pathological classification models
HisImage [42] (hard vote) 75.31 69.36 65.60 69.26
FMDNN [26] (hard vote) 92.59 93.73 90.17 91.37
StoHisNet [43] (hard vote) 85.23 75.84 67.09 84.23
HiFuse [33] (hard vote) 64.20 43.76 47.43 43.80
Ours
CMSwinKAN-tiny (hard vote) 95.06 96.69 93.40 95.34
We conducted WSI classification experiments on the publicly available breast cancer histopathological dataset, BreakHis, to further confirm its generalization ability. As shown in Table 7, when using a hard voting strategy, CMSwinKAN exhibited outstanding performance even in imbalanced data scenarios: its ACC (95.06%), BACC (96.69%), KAPPA (93.40%), and F1 (95.34%) significantly surpassed all comparative models. Compared to the current optimal large pathological model, CONCH (ACC 95.05%), CMSwinKAN won by a narrow margin and led considerably by 2.94% in the BACC metric (93.75% to 96.69%); its performance improved by 1.23% compared to the classic architecture DenseNet (ACC 93.83%). This result is consistent with the conclusions drawn from the PpNTs dataset, confirming that the model effectively adapts to pathological image classification needs across different data distributions through multi-scale feature interaction with the CMSA module and nonlinear mapping in the KAN layer.
4.6 Ablation Experiments
4.6.1 Model Performance Ablation
Our ablation experiments further analyzed the impact of the KAN network, CDFA component, and CMSA on classification performance across both datasets. We assessed the performance of the CMSwinKAN model after restoring the KAN network, replacing the CDFA component with a linear layer, removing the CMSA module, and comparing it to the complete CMSwinKAN model. Table 8 confirms that these three modules positively influence the model’s classification performance. Specifically:
Impact of the KAN Network: In the SwinKANsformer Block, the KAN network replaces the traditional MLP layer, enhancing the model’s ability to extract nonlinear features and improving interpretability. The conventional linear layer classification head is replaced with a more efficient KAN network. If the KAN network is restored and the MLP layer and linear classification head are used instead, the classification performance significantly declines. Specifically, ACC drops from 93.58% to 92.29%, BACC from 93.42% to 92.09%, and KAPPA and F1 scores decrease from 92.12% and 93.57% to 90.93% and 92.28%, respectively. This decline is due to a significant reduction in feature expression capability when reverting to the MLP layer and linear classification head, which fails to adequately extract fine-grained features in pathological slice images, leading to overly similar learned features that easily confuse different cell types.
Replacing the CDFA Module with a Linear Layer: The CDFA module enhances the model’s learning ability for information from global to microscopic scales through feature comparison across different scales. When the CDFA module is replaced with a linear layer, effective interaction between different scales is lost, resulting in significant feature information loss and a marked decline in all metrics. Specifically, KAPPA decreases by 1.29%, and F1 scores and other metrics drop by 0.7%. This highlights the crucial role of the CDFA module in multi-scale feature fusion, significantly aiding the model’s ability to integrate high-level semantic information and detailed edge information in pathological images.
Table 8: Ablation studies of CMSwinKAN on our private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC(%)
w/o KAN 92.29 92.09 90.93 92.28 99.39
w/o CDFA 92.89 92.73 90.93 92.88 99.44
w/o CMSA 92.50 92.24 90.85 92.49 99.42
CMSwinKAN 93.58 93.42 92.12 93.57 99.53
Table 9: Ablation studies of CMSwinKAN on the public BreakHis dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC(%)
w/o KAN 80.66 79.59 77.31 80.90 97.81
w/o CDFA 82.18 82.41 75.75 82.60 98.14
w/o CMSA 79.64 81.40 77.44 80.06 97.78
CMSwinKAN 87.79 87.56 85.69 88.04 98.80
Removing the CMSA Module: The CMSA module combines the CDFA module with a learnable weight allocator, adjusting the contribution distribution of multi-scale features to capture long-range dependencies and local features in pathological slice images while avoiding contamination from scales with excessive low-density information. When the CMSA module is removed, the complete loss of multi-scale details leads to a significant decline in the model’s ability to capture global and local features, resulting in noticeable reductions in performance metrics. Specifically, ACC decreased by 1.15%, BACC by 1.28%, and KAPPA by 1.38%.
Notably, the performance decline from removing the CMSA module is more pronounced than replacing the CDFA module with a linear layer, reflecting the effectiveness of the CMSA module. The multi-scale contribution weight allocation significantly supports the model in extracting effective high-density feature information, thereby enhancing the powerful capabilities of multi-scale feature learning when combined with the CDFA module.
4.6.2 Validation of Model Generalization Ability
Additionally, we conducted the same ablation experiments on the publicly available breast tissue histopathological image classification dataset, BreakHis, which features an imbalanced class distribution, further reflecting the impact of different components on model classification performance. These experiments further validate the significant influence of each key component on classification accuracy. As shown in Table 9, removing any component leads to a marked decline in classification metrics, highlighting the importance of these components in feature extraction and expression capabilities, as well as the necessity of effectively utilizing high-level semantics and detailed edges in pathological images to simulate clinical pathology diagnostics, thus enhancing the model’s classification ability.
5 Conclusion
In this paper, we proposed a model named CMSwinKAN for the patch classification task of NB pathological slices stained with H&E. This model addressed several issues present in existing methods, such as insufficient integration of global features and local edge details, poor model interpretability, and the susceptibility of multi-scale features at different stages to contamination from low-density information. Additionally, we designed a soft voting mechanism based on prior knowledge of tissue composition to facilitate an effective transition from patch-level classification to WSI-level classification.
We evaluated the performance of the proposed CMSwinKAN model against other classical and advanced methods for both patch-level and WSI-level classification on the private PpNTs dataset and the publicly available BreakHis dataset. The experimental results demonstrated that CMSwinKAN significantly outperforms existing pathological image classification methods on these datasets and markedly enhances model performance in WSI classification tasks through the soft voting mechanism, achieving excellent results across nearly all evaluation metrics. Furthermore, we conducted ablation experiments to validate the importance of key components within the model, and the results confirmed that these components contributed positively and significantly to model performance.